40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 16 10:01:16 2010, Alberto, Configuration, LSC, Arms and PRC not locking 
    Reply  Tue Feb 16 14:10:06 2010, josephb, rob, koji, Configuration, LSC, Arms and PRC not locking 
       Reply  Tue Feb 16 15:25:00 2010, Alberto, Configuration, LSC, Arms and PRC not locking shock.jpg
Message ID: 2605     Entry time: Tue Feb 16 10:01:16 2010     Reply to this: 2607
Author: Alberto 
Type: Configuration 
Category: LSC 
Subject: Arms and PRC not locking 

Since last Friday either the arms or the PRC can't lock.

The montors show the beam flashing on the end mirrors, but the cavity can't get locked. The error signal looks fine. I suspect a computer problem.

Also PRC can't lock. SPOB is suspiciously stuck at about -95. Although that's not a fixed number, but covering the by hand the SPOB PD on the ITMY table doesn't change the number. I check the DC output of the photodetector and it is actually seen the beam.

Suspecting computer problems started after last Thursday's IP switch, I rebooted the frame builder, c1dcuepics, c1daqctrl and all the front ends. I then burtrestored to February 1st at 1:00 am.

Before I burtrestored c1iscepics, SPOB had gone back to more typical numbers around 0, as it usually read when PRC wasn't locked.

But burtrestoring c1iscepics, return it to the -95 of earlier.

Burterestoring to other times or dates didn't solve the problems.

ELOG V3.1.3-